p/3 - 3p/8 = 4
I have no idea where to start :P

Answer by nerdybill(7384)

p/3 - 3p/8 = 4
Multiply both sides by 24 to get rid of denominators:
8p - 9p = 96
-p = 96
p = -96
.
To check, plug back into original equation:
p/3 - 3p/8 = 4
-96/3 - 3(-96)/8 = 4
-32 - 3(-12) = 4
-32 + 36 = 4
4 = 4